AMERICAN UNIVERSITY Calculate 314 (s+3) (4s+5) 1 A- B . I 00 c?
Answer #1

we have

L^{-1}\left [ \frac{s+4}{(s+3)(s^2+4s+5)} \right ]

by the partial fraction,

s+4 (s +3)(52 + 4s + 5) = A (s +3) + Bs + sa + 4s + 5)

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{A(s^2+4s+5)+(Bs+C)(s+3)}{(s+3)(s^2+4s+5)}

s+4=A(s^2+4s+5)+(Bs+C)(s+3) ....................1)

put s = -3 in equation 1),

-3+4=A((-3)^2+4(-3)+5)+(Bs+C)(-3+3)

1=A(2)+0

A=\frac{1}{2}

put A = 1/2 in equation 1),

s+4=\frac{1}{2}(s^2+4s+5)+Bs^2+3Bs+Cs+3C

s+4-\frac{1}{2}s^2-2s-\frac{5}{2}=Bs^2+3Bs+Cs+3C

-\frac{1}{2}s^2-s+\frac{3}{2}=Bs^2+(3B+C)s+3C

compare both side,

B=-\frac{1}{2}, \ \ 3C=\frac{3}{2}

B=-\frac{1}{2}, \ \ C=\frac{1}{2}

so,

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{\frac{1}{2}}{(s+3)}+\frac{-\frac{1}{2}s+\frac{1}{2}}{(s^2+4s+5)}

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{1}{2(s+3)}+\frac{1}{2}\left [\frac{-s+1}{s^2+4s+4+1} \right ]

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{1}{2(s+3)}+\frac{1}{2}\left [\frac{-s-2+3}{(s+2)^2+(1)^2} \right ]

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{1}{2(s+3)}+\frac{1}{2}\left [-\frac{(s+2)}{(s+2)^2+(1)^2}+\frac{3}{(s+2)^2+(1)^2} \right ]

\frac{s+4}{(s+3)(s^2+4s+5)}=\frac{1}{2}\left [\frac{1}{(s+3)} \right ]-\frac{1}{2}\left [\frac{(s+2)}{(s+2)^2+(1)^2} \right ]+\frac{3}{2}\left [\frac{1}{(s+2)^2+(1)^2} \right ]

now,

L^{-1}\left [\frac{s+4}{(s+3)(s^2+4s+5)} \right ]=\frac{1}{2}L^{-1}\left [\frac{1}{(s+3)} \right ]-\frac{1}{2}L^{-1}\left [\frac{(s+2)}{(s+2)^2+(1)^2} \right ]+\frac{3}{2}L^{-1}\left [\frac{1}{(s+2)^2+(1)^2} \right ]

L^{-1}\left [\frac{s+4}{(s+3)(s^2+4s+5)} \right ]=\frac{1}{2}e^{-3t}-\frac{1}{2}e^{-2t}cos(t)+\frac{3}{2}e^{-2t}sin(t)